OpenStudy (anonymous):

Kelly makes $10 more per hour than her regular pay when she is working overtime. Yesterday, she worked 8 regular hours and 4 overtime hours and earned $160. What is Kelly's regular pay per hour?

Directrix (directrix):

What is Kelly's regular pay per hour? (A) $20 per hour (B) $10 per hour (C) $12 per hour (D) $16 per hour

Directrix (directrix):

Let x = K's pay per regular hour Then, K's pay per overtime hour is (x + 10) 8x + 4(x + 10) = 160 8x + 4x + 40 = 160 Solve for x to get K's pay per regular hour @StudiousHojea

OpenStudy (annie96):

(B) $10 per hour 8x+4(x+10)=160 x=10 dollars